Using identities, evaluate.
 
998²
 
Please answer fast it is my holiday homework

Grade:8

1 Answers

Ayush 8A 7239
2 years ago
9982 = (1000-2)2
using (a-b) 2 = a2 + b2 – 2ab
(1000-2)2 = 1000000 + 4 – 2 × 1000 × 2
                = 1000000 + 4 – 4000
                = 996004